OpenStudy (anonymous):

The equation y = 16x + 560 describes the number of calories in a recipe, where y is the total number of calories and x is the number of teaspoons of sugar added. The table shows the number of calories in another recipe, where y is the total number of calories and x is the number of teaspoons of honey added. How many calories are in a teaspoon of sugar, and how many are in a teaspoon of honey? Teaspoons of honey 1 2 4 7 Total calories 504 568 696 888 A teaspoon of sugar has ____ calories. A teaspoon of honey has ____ calories.

OpenStudy (anonymous):

@pooja195 @Qwertty123 @MissSmartiez @myininaya

OpenStudy (jdoe0001):

so ahemm take a peek at your table for the calories in 1 teaspoon of honey so \(\begin{array}{rcccc} honey&1&2&4&7\\ calories&504&568&696&888 \end{array}\) how many do you think?

OpenStudy (anonymous):

well for honey i know theres 504 calories

OpenStudy (anonymous):

i don't know about sugar

OpenStudy (jdoe0001):

ok now for the sugar added recall "sugar added" = x "calories" = y so how many "y" n 1 teaspoon of sugar, or in 1 x or if x = 1, what's "y"? \(\bf y=16x+560\qquad {\color{brown}{ x=1}}\implies y=16{\color{brown}{ (1)}}+560\)

OpenStudy (anonymous):

so sugar calories would be 577

OpenStudy (jdoe0001):

577?

OpenStudy (anonymous):

577 calories

OpenStudy (jdoe0001):

well.. hmm how did you get 577 though?

OpenStudy (anonymous):

16+1+560=577

OpenStudy (anonymous):

wait

OpenStudy (jdoe0001):

\(\bf y=16x+560\qquad {\color{brown}{ x=1}}\implies y=16{\color{brown}{ (1)}}+560\implies y=16\cdot 1+560\)

OpenStudy (anonymous):

so 576?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thank you soooooo much :)

OpenStudy (jdoe0001):

yw